1. A dealer provides a customer with the following American term quotes on the British pound:

Bid Ask
Pound 1.2391 1.2396

a. Is the pound the asset or the price in the quotes provided by the dealer? Circle one.

ASSET PRICE

b. Use the information provided to calculate the dealer’s European term quotes on the pound. Enter the bid and ask values (4 decimal places) in the table below.For full credit you must show your work.

Bid Ask
Pound

2. Philip has just completed transactions in France and Japan. He is now holding €100,000 and ¥300,000 and wants to convert to U.S. dollars. His currency dealer provides this quotation:

USD/JPY: 0.010094 – 97

EUR/USD: 0.7638 – 40

What are Philip’s proceeds from conversion? For full credit you must show your work
Answer #1

1]

a]

In an American term quote, the domestic currency (dollar) is the quote currency, and the foreign currency (pound) is the base currency.

Hence, in this quote, pound is the asset

b]

pound bid = 1 / American term quote ask = 1 / 1.2396 = 0.8067

pound ask = 1 / American term quote bid = 1 / 1.2391 = 0.8070